Convert 53 cm/sec to in/min. 1 in = 2.54 cm, 1min = 60 sec.
Answer by Alan3354(69443)

Convert 53 cm/sec to in/min. 1 in = 2.54 cm, 1min = 60 sec.
----------------------
53 cm = 53 cm*(1 in/2.54 cm) = 53/2.54 inches
Notice the cms cancel
----
53 cm/sec = 53/2.54 in/sec * 60 sec/1 min = (53/2.54)*60
Notice the seconds cancel
= ~1252 in/minute
====================
Including the units reduces the chance of multiplying when you should divide and v.v.